### v1

@@
Acknowledgements Part of this research was carried out at the Jet Propulsion Laboratory, California Institute of Technology, under a contract with the National Aeronautics and Space Administration. Basic research in radio astronomy at the Naval Research Laboratory is supported by 6.1 Base funding. The NANOGrav Collaboration receives support from National Science Foundation Physics Frontiers Center award numbers 1430284 and 2020265. The National Radio Astronomy Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc. A.C. acknowledges support from the National Science Foundation, the National Aeronautics and Space Administration, and the U.S. Department of Energy.
This report reflects an initial set of recommendations by the Transition Advisory Group for the next-generation Very Large Array Project and is distributed for the purposes of obtaining community comment. Modification of this report in response to community comment is expected.
@@
In order to assess risks and benefits of possible transition plans in between the above two extremes, the ngVLA project established the VLA+VLBA to ngVLA Transition Advisory Group (TAG)—a group of 18 members of the U.S. and international astronomical community (including members of the ngVLA Science Advisory Council). The TAG was charged with identifying (i) scientific opportunities in the coming decade that will critically benefit from complementary and/or unique observations at radio wavelengths; (ii) key stakeholders for the transition, (iii) parameters of interest by which to compare transition options; and (iv) necessary metrics by which to quantify the impact of different transition options. The primary deliverable from the TAG is a “VLA+VLBA to ngVLA Transition Option Concepts” report that includes a prioritized list of transition options.
The TAG assembled a set of VLA-VLBA-ngVLA science use cases, leveraging a variety of resources including the ngVLA Science Book, NRAO press releases, and, ultimately, the Pathways to Discovery Decadal Survey. The science use cases were augmented by discussion of the VLA in the Origins, Worlds, and Life (Planetary Science & Astrobiology) Decadal Survey and the Exploring and Safeguarding Humanity’s Home in Space (Solar & Space Physics) Decadal Survey. These science cases, and particularly those identified as ones to which the VLA or VLBA or both would make a “very significant” contribution (as judged by Pathways to Discovery), were used to assess the scientific impact of a number of technical options provided by an NRAO Internal Technical Analysis Team (ITAT). The NRAO also provided a nominal plan for the design, development, and construction of the ngVLA, which envisions a 10 year construction interval with early science beginning three years after the start of construction.
@@ 2 Charge and Evaluation Process
While conducting its work, the TAG advertised its activities broadly and received input from the scientific community (§§78). The TAG plans to make this report available in a public forum, e.g., arXiv, and accept comments for up to 90 days before producing a final version.
@@ 4 Technical Options for the Transition
The TAG received from the NRAO ITAT a document entitled “Transition Supporting Information” outlining a set of technical options that could be considered in implementing the VLA+VLBA to ngVLA transition. In general, these technical options involve reducing some measure of capability for either the VLA or the VLBA or both, with the aim of enabling NRAO to redirect staff efforts from VLA or VLBA operations to ngVLA construction. The TAG consolidated these technical options into the following broad set, listed in no particular order:
@@ 6.3 Second VLA Transition Option
Should the first transition option be infeasible or insufficient, the recommended second option is to reduce the set of receivers at each antenna, provided that at least five of the current frequency bands are maintained at all antennas. The TAG judges that no fewer than five frequency bands are required in order to enable the VLA to have a sufficient spectral dynamic range (§5). Further, at a minimum, these five frequency bands must span from “low” frequency to “high” frequency. Here, “low” frequency is considered to be C band (4 GHz–8 GHz) and lower frequencies while “high” frequency is considered to be Ku band (12 GHz–18 GHz) and higher frequencies. X band (8 GHz–12 GHz) must be maintained at all times as it is required for reference pointing calibration for higher frequencies, and X band is one of the most heavily-demanded bands on the VLA. Based on these considerations, the baseline set of receivers for the transition is L-S-C-X-K.
@@ 6.4 Third VLA Transition Option
Given that no single arrangement of antennas appears capable of retaining the VLA’s combination of surface brightness sensitivity and angular resolution capabilities (ngvla126, ), a fixed configuration should be adopted only if necessary.
@@ 8 Stakeholder Engagement
  • A Splinter Session entitled “The Next-Generation VLA: Update and Community Forum” was organized at the $245^{\mathrm{th}}$ Meeting of the American Astronomical Society (2025 January) in order to present an overview of the current status and seek community input.
In addition to these activities, the TAG plans to conduct a post the initial report on arXiv (astro-ph) in order to seek comments, with at least 90 days provided for community members to submit comments, and to conduct a Webinar. These comments will be assessed as part of developing the final Transition report.

### v2

@@
Acknowledgements A.C. acknowledges support from the National Science Foundation, the National Aeronautics and Space Administration, and the U.S. Department of Energy. Part of this research was carried out at the Jet Propulsion Laboratory, California Institute of Technology, under a contract with the National Aeronautics and Space Administration. L.L. acknowledges the support of UNAM-DGAPA PAPIIT grants IN112820 and IN108324, and CONACYT-CF grant 263356. Basic research in radio astronomy at the Naval Research Laboratory is supported by 6.1 Base funding. The NANOGrav Collaboration receives support from National Science Foundation Physics Frontiers Center award numbers 1430284 and 2020265. The National Radio Astronomy Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc.
@@
In order to assess risks and benefits of possible transition plans in between the above two extremes, the ngVLA project established the VLA+VLBA to ngVLA Transition Advisory Group (TAG)—a group of 18 members of the U.S. and international astronomical community (including members of the ngVLA Science Advisory Council). The TAG was charged with identifying (i) scientific opportunities in the coming decade that will critically benefit from complementary and/or unique observations at radio wavelengths; (ii) key stakeholders for the transition; (iii) parameters of interest by which to compare transition options; and (iv) necessary metrics by which to quantify the impact of different transition options. The primary deliverable from the TAG is a “VLA+VLBA to ngVLA Transition Option Concepts” report that includes a prioritized list of transition options.
The TAG assembled a set of VLA-VLBA-ngVLA science use cases, leveraging a variety of resources including the ngVLA Science Book, NRAO press releases, and, ultimately, the Pathways to Discovery Decadal Survey. The science use cases were augmented by discussion of the VLA in the Origins, Worlds, and Life (Planetary Science & Astrobiology) Decadal Survey and the Exploring and Safeguarding Humanity’s Home in Space (Solar & Space Physics) Decadal Survey. These science cases, and particularly those identified as ones to which the VLA or VLBA or both would make a “very significant” contribution (as judged by Pathways to Discovery), were used to assess the scientific impact of a number of technical options provided by an NRAO Internal Technical Analysis Team (ITAT). Considerations of transition options were limited to hardware, and did not encompass other critical aspects of the VLA or VLBA such as the implementation of scientific operations, data processing, and data products delivery to the community. The NRAO also provided a nominal plan for the design, development, and construction of the ngVLA, which envisions a 10 year construction interval with early science beginning three years after the start of construction.
@@ 2 Charge and Evaluation Process
While conducting its work, the TAG advertised its activities broadly and received input from the scientific community (§§78). The TAG made this report available in a public forum (arXiv) and accept comments for up to 90 days before producing this final version.
@@ 4 Technical Options for the Transition
The TAG received from the NRAO ITAT a document entitled “Transition Supporting Information” outlining a set of technical options that could be considered in implementing the VLA+VLBA to ngVLA transition. In general, these technical options involve reducing some measure of capability for either the VLA or the VLBA or both, with the aim of enabling NRAO to redirect staff efforts from VLA or VLBA operations to ngVLA construction. Further, the transition options were limited to hardware capabilities and did not encompass other critical aspects of the VLA and VLBA, such as the implementation of scientific operations, data processing, and data products delivery to the community.
The TAG consolidated these technical options into the following broad set, listed in no particular order:
@@
There also likely are a number of intermediate or “hybrid” approaches. For instance, rather than retain the current full VLA receiver suite or shut down some of the receivers, the VLA might be able to be used as a testbed for ngVLA-like receivers. There have been multiple designs published for wideband receivers (e.g., 8879392, ; 2023IJMw….3..570K, ), and one of these designs could be tested on the VLA antennas, allowing NRAO staff to begin to transition to maintaining ngVLA-like equipment while maintaining essentially full VLA capabilities. Alternately, there could be an “early deployment” of ngVLA correlation or data processing, shifting such capabilities rapidly from the current VLA or VLBA to the ngVLA. This latter approach may reduce the risk that current VLA or VLBA capabilities are increasingly difficult to maintain due to obsolescence of digital signal processing or computational equipment. The TAG did not consider such possibilities at length.
@@ 6.3 Second VLA Transition Option
Should the first transition option be infeasible or insufficient, the recommended second option is to reduce the set of receivers at each antenna, provided that at least five of the current frequency bands are maintained at all antennas. The TAG judges that no fewer than five frequency bands are required in order to enable the VLA to have a sufficient spectral dynamic range (§5). Further, at a minimum, these five frequency bands must span from “low” frequency to “high” frequency. Here, “low” frequency is considered to be C band (4 GHz–8 GHz) and lower frequencies while “high” frequency is considered to be Ku band (12 GHz–18 GHz) and higher frequencies. X band (8 GHz–12 GHz) must be maintained at all times as it is required for reference pointing calibration for higher frequencies, and X band is one of the most heavily-demanded bands on the VLA. Based on these considerations, the baseline set of receivers for the transition is L-S-C-X-K. We note that this suite of receivers does not include either the low-frequency (P-band) or Q-band receivers, which has the effect of reducing the extent to which the VLA’s frequency range is unique.
@@ 6.4 Third VLA Transition Option
Given that no single arrangement of antennas appears capable of retaining the VLA’s combination of surface brightness sensitivity and angular resolution capabilities (ngvla126, ), and the potential for increased difficulties in maintaining the antennas (§5.2.2), a fixed configuration should be adopted only if necessary.
@@ 7 Other Considerations
While the TAG recommends that the Transition be of as limited duration as possible (§6.2), there is the possibility that VLA observations would have to cease for some interval. Even so, the VLA Archive remains a rich source of scientific results and a means by which early-career researchers can gain experience with radio astronomical interferometers. The TAG did not discuss such possibilities at length, but a limited-duration “archival data analysis” program may be warranted to ensure continued engagement between the current NRAO user community and NRAO facilities, should there have to be a cessation of new VLA observations.
@@ 8 Stakeholder Engagement
  • A Splinter Session entitled “The Next-Generation VLA: Update and Community Forum” was organized at the $245^{\mathrm{th}}$ Meeting of the American Astronomical Society (2025 January) in order to present an overview of the current status and seek community input. The TAG has posted its initial report on arXiv (astro-ph) for a 90-day comment period for community members to submit comments. These comments have been addressed in this final version of the report.